### v1

# Generalization in Reinforcement Learning for Radio Access Networks

Abstract
[…]
AI
artificial intelligence
ACK
positive acknowledgment
ARQ
automatic repeat request
BE
baseline encoding
BC
behavioral cloning
BCQ
batch-constrained deep Q-learning
CCTR
Channel-Conditioned Target Return
CNN
convolutional neural network
CPU
central processing unit
CQL
conservative Q-learning
CT
continuous time
CV
computer vision
DAVG
discounted average
DLLA
downlink link adaptation
DME
data management entity
DP
dynamic programming
DQN
deep Q-network
DT
decision transformer
eMBB
enhanced mobile broadband
FB
full buffer
ILLA
inner-loop link adaptation
gNB
next generation NodeB
gNB-CU
gNB centralized unit
gNB-DU
gNB distributed unit
GAT
graph attention network
GNN
graph neural networks
GCN
graph convolutional network
GPI
generalized policy iteration
GPU
graphics processing unit
HPC
high-performance computing
LA
link adaptation
LN
layer normalization
LCM
life cycle management
LSF
load sharing facility
LSTM
long short-term memory
LT
learnable time
MAB
multi-armed bandits
MC
Monte Carlo
MDP
Markov decision process
ML
machine learning
MLP
multi-layer perceptron
mMIMO
massive MIMO
mMTC
massive machine-type communications
NACK
negative acknowledgment
NF
network function
NLP
natural language processing
OLLA
outer-loop link adaptation
ORAN
Open RAN
PE
positional encoding
PHY
physical layer
RL
reinforcement learning
RLC
radio link control
RNN
recurrent neural network
RTGs
returns-to-go
RTG
return-to-go
RT RIC
nea-real-time RAN intelligent controller
RvS
reinforcement learning via supervised learning
SACo
state-action coverage
SADCo
state-action density coverage
SE
spectral efficiency
SMO
service management and orchestration
SMOF
SMO function
TBS
transport block size
TD
temporal difference
TQ
relative trajectory quality
URLLC
ultra-reliable low-latency communications
ZSG
zero-shot generalization
VAE
variational auto-encoder
2G
Second Generation
3G
[…]
3GPP
[…]
4G
[…]
5G
[…]
6G
[…]
AA
Antenna Array
AC
Admission Control
AD
Attack-Decay
ADSL
Asymmetric Digital Subscriber Line
AHW
Alternate Hop-and-Wait
AMC
Adaptive Modulation and Coding
AoA
angle of arrival
AP
Access Point
APA
Adaptive Power Allocation
AR
autoregressive
ARMA
Autoregressive Moving Average
ATES
Adaptive Throughput-based Efficiency-Satisfaction Trade-Off
AWGN
additive white Gaussian noise
BB
Branch and Bound
BD
Block Diagonalization
BER
bit error rate
BF
Best Fit
BLER
block error rate
BPC
Binary power control
BPSK
Binary Phase-Shift Keying
BPA
Best pilot-to-data power ratio (PDPR) Algorithm
BRA
Balanced Random Allocation
BCRB
Bayesian Cramér-Rao Bound
BS
base station
CAP
Combinatorial Allocation Problem
CAPEX
Capital Expenditure
CBF
Coordinated Beamforming
CBR
Constant Bit Rate
CBS
Class Based Scheduling
CC
Congestion Control
CDF
cumulative distribution function
CDMA
Code-Division Multiple Access
CL
Closed Loop
CLPC
Closed Loop Power Control
CNR
Channel-to-Noise Ratio
CPA
Cellular Protection Algorithm
CPICH
Common Pilot Channel
CoMP
Coordinated Multi-Point
CQI
channel quality indicator
CRB
Cramér-Rao Bound
CRM
Constrained Rate Maximization
CRN
Cognitive Radio Network
CS
Coordinated Scheduling
CSI
channel state information
CSIR
channel state information at the receiver
CSIT
channel state information at the transmitter
CUE
cellular user equipment
D2D
device-to-device
DCA
Dynamic Channel Allocation
DE
Differential Evolution
DFT
Discrete Fourier Transform
DIST
Distance
DL
downlink
DMA
Double Moving Average
DMRS
demodulation reference signal
D2DM
D2D Mode
DMS
D2D Mode Selection
DPC
Dirty Paper Coding
DRA
Dynamic Resource Assignment
DSA
Dynamic Spectrum Access
DSM
Delay-based Satisfaction Maximization
ECC
Electronic Communications Committee
EFLC
Error Feedback Based Load Control
EI
Efficiency Indicator
eNB
Evolved Node B
EPA
Equal Power Allocation
EPC
Evolved Packet Core
EPS
Evolved Packet System
ESPRIT
estimation of signal parameters via rotational invariance
E-UTRAN
Evolved Universal Terrestrial Radio Access Network
ES
Exhaustive Search
FDD
frequency division duplexing
FDM
Frequency Division Multiplexing
FER
Frame Erasure Rate
FF
Fast Fading
FIM
Fisher information matrix
FSB
Fixed Switched Beamforming
FST
Fixed SNR Target
FTP
File Transfer Protocol
GA
Genetic Algorithm
GBR
Guaranteed Bit Rate
GLR
Gain to Leakage Ratio
GOS
Generated Orthogonal Sequence
GPL
GNU General Public License
GRP
Grouping
HARQ
hybrid automatic repeat request
HMS
Harmonic Mode Selection
HOL
Head Of Line
HSDPA
High-Speed Downlink Packet Access
HSPA
High Speed Packet Access
HTTP
HyperText Transfer Protocol
ICMP
Internet Control Message Protocol
ICI
Intercell Interference
ID
Identification
ISAC
integrated sensing and communication
IEEE
Institute of Electrical and Electronics Engineers
IETF
Internet Engineering Task Force
ILP
Integer Linear Program
JRAPAP
Joint RB Assignment and Power Allocation Problem
UID
Unique Identification
IID
Independent and Identically Distributed
IIR
Infinite Impulse Response
ILP
Integer Linear Problem
IMT
International Mobile Telecommunications
INV
Inverted Norm-based Grouping
IoT
Internet of Things
IP
Internet Protocol
IPv6
Internet Protocol Version 6
ISD
Inter-Site Distance
ISI
Inter Symbol Interference
ITU
International Telecommunication Union
JOAS
Joint Opportunistic Assignment and Scheduling
JOS
Joint Opportunistic Scheduling
JP
Joint Processing
JS
Jump-Stay
KKT
Karush-Kuhn-Tucker
L3
Layer-3
LAC
Link Admission Control
LC
Load Control
LOS
Line of Sight
LP
Linear Programming
LS
least squares
LTE
Long Term Evolution
LTE-A
LTE-Advanced
LTE-Advanced
Long Term Evolution Advanced
M2M
Machine-to-Machine
MAC
Medium Access Control
MANET
Mobile Ad hoc Network
MCS
modulation and coding scheme
MDB
Measured Delay Based
MDI
Minimum D2D Interference
MF
Matched Filter
MG
Maximum Gain
MH
Multi-Hop
MIMO
multiple input multiple output
MINLP
Mixed Integer Nonlinear Programming
MIP
Mixed Integer Programming
MISO
Multiple Input Single Output
MLE
maximum likelihood estimator
MLWDF
Modified Largest Weighted Delay First
MME
Mobility Management Entity
MMSE
minimum mean squared error
MOS
Mean Opinion Score
MPF
Multicarrier Proportional Fair
MRA
Maximum Rate Allocation
MR
Maximum Rate
MRC
Maximum Ratio Combining
MRT
Maximum Ratio Transmission
MRUS
Maximum Rate with User Satisfaction
MS
mobile station
MSE
mean squared error
MSI
Multi-Stream Interference
MTC
Machine-Type Communication
MTSI
Multimedia Telephony Services over IMS
MTSM
Modified Throughput-based Satisfaction Maximization
MU-MIMO
multiuser multiple input multiple output
MU
multi-user
MUSIC
multiple signal classification
NAS
Non-Access Stratum
NB
Node B
NE
Nash equilibrium
NCL
Neighbor Cell List
NLOS
Non-Line of Sight
NMSE
Normalized Mean Square Error
NORM
Normalized Projection-based Grouping
NP
Non-Polynomial Time
NR
New Radio
NRT
Non-Real Time
NSPS
National Security and Public Safety Services
O2I
Outdoor to Indoor
OFDMA
orthogonal frequency division multiple access
OFDM
orthogonal frequency division multiplexing
OFPC
Open Loop with Fractional Path Loss Compensation
O2I
Outdoor-to-Indoor
OL
Open Loop
OLPC
Open-Loop Power Control
OL-PC
Open-Loop Power Control
OPEX
Operational Expenditure
ORB
Orthogonal Random Beamforming
JO-PF
Joint Opportunistic Proportional Fair
OSI
Open Systems Interconnection
PAIR
D2D Pair Gain-based Grouping
PAPR
Peak-to-Average Power Ratio
P2P
Peer-to-Peer
PC
Power Control
PCI
Physical Cell ID
PDF
Probability Density Function
PDPR
pilot-to-data power ratio
PER
Packet Error Rate
PF
Proportional Fair
P-GW
Packet Data Network Gateway
PL
Pathloss
PPR
pilot power ratio
PRB
physical resource block
PROJ
Projection-based Grouping
ProSe
Proximity Services
PS
Packet Scheduling
PSAM
pilot symbol assisted modulation
PSO
Particle Swarm Optimization
PZF
Projected Zero-Forcing
QAM
Quadrature Amplitude Modulation
QoS
Quality of Service
QPSK
Quadri-Phase Shift Keying
RAISES
Reallocation-based Assignment for Improved Spectral Efficiency and Satisfaction
RAN
radio access network
RAT
Radio Access Technology
RATE
Rate-based
RB
resource block
RBG
Resource block broup
REF
Reference Grouping
RM
Rate Maximization
RNC
Radio Network Controller
RND
Random Grouping
RRA
Radio Resource Allocation
RRM
radio resource management
RSCP
Received Signal Code Power
RSRP
Reference Signal Receive Power
RSRQ
Reference Signal Receive Quality
RR
Round Robin
RRC
Radio Resource Control
RSSI
Received Signal Strength Indicator
RT
Real Time
RU
Resource Unit
RUNE
RUdimentary Network Emulator
RV
Random Variable
SAC
Session Admission Control
SCM
Spatial Channel Model
SC-FDMA
Single Carrier - Frequency Division Multiple Access
SD
Soft Dropping
S-D
Source-Destination
SDPC
Soft Dropping Power Control
SDMA
Space-Division Multiple Access
SER
Symbol Error Rate
SES
Simple Exponential Smoothing
S-GW
Serving Gateway
SINR
signal-to-interference-plus-noise ratio
SI
Satisfaction Indicator
SIP
Session Initiation Protocol
SISO
single input single output
SIMO
Single Input Multiple Output
SIR
signal-to-interference ratio
SLNR
Signal-to-Leakage-plus-Noise Ratio
SMA
Simple Moving Average
SNR
signal-to-noise ratio
SORA
Satisfaction Oriented Resource Allocation
SORA-NRT
Satisfaction-Oriented Resource Allocation for Non-Real Time Services
SORA-RT
Satisfaction-Oriented Resource Allocation for Real Time Services
SPF
Single-Carrier Proportional Fair
SRA
Sequential Removal Algorithm
SRS
Sounding Reference Signal
SSB
synchronisation signal block
SU-MIMO
single-user multiple input multiple output
SU
Single-User
SVD
Singular Value Decomposition
TCP
Transmission Control Protocol
TDD
time division duplexing
TDMA
Time Division Multiple Access
TETRA
Terrestrial Trunked Radio
TP
Transmit Power
TPC
Transmit Power Control
TTI
transmission time interval
TTR
Time-To-Rendezvous
TSM
Throughput-based Satisfaction Maximization
TU
Typical Urban
UE
user equipment
UEPS
Urgency and Efficiency-based Packet Scheduling
UL
uplink
UMTS
Universal Mobile Telecommunications System
URI
Uniform Resource Identifier
URM
Unconstrained Rate Maximization
UT
user terminal
VR
Virtual Resource
VoIP
Voice over IP
WAN
Wireless Access Network
WCDMA
Wideband Code Division Multiple Access
WF
Water-filling
WiMAX
Worldwide Interoperability for Microwave Access
WINNER
Wireless World Initiative New Radio
WLAN
Wireless Local Area Network
WMPF
Weighted Multicarrier Proportional Fair
WPF
Weighted Proportional Fair
WSN
Wireless Sensor Network
WWW
World Wide Web
XIXO
(Single or Multiple) Input (Single or Multiple) Output
ZF
zero-forcing
ZMCSCG
Zero Mean Circularly Symmetric Complex Gaussian
1 INTRODUCTION
Wireless communication networks have rapidly evolved into increasingly complex and heterogeneous systems, underscoring the need for advanced automation and optimization. Managing modern RANs, such as the 5G New Radio (NR) system, requires solving large-scale, often combinatorial problems under strict latency constraints—particularly those encountered in RRM. The highly dynamic and stochastic nature of the RAN environment further challenges the effectiveness of classical rule-based heuristics, which struggle to adapt to rapidly changing network conditions and diverse service requirements. These limitations have prompted a shift toward adaptive, data-driven solutions.
[…]
However, several studies Farebrother et al. (2018); Zhang et al. (2018b, a); Gamrian and Goldberg (2019); Song et al. (2019); Igl et al. (2019) have highlighted the limited generalization ability of RL agents to transfer learned knowledge to novel environments. This limitation primarily arises from training algorithms within fixed environments, which results in overfitting to specific transition dynamics and reward structures Farebrother et al. (2018). Consequently, such models perform well under in-distribution conditions but suffer significant degradation when exposed to unseen environments. The challenge of RL generalization is exacerbated in partially observable or uncertain Markov decision processes, where incomplete knowledge or environmental uncertainty renders available states insufficient for optimal decision-making Igl et al. (2019).
To generalize effectively, RL agents must learn transferable representations that are robust to variations in the environment, including uncertainty and feature variability Kirk et al. (2023). While classical regularization techniques from supervised learning—such as weight decay, dropout, and batch or layer normalization—can enhance the robustness of RL policies Cobbe et al. (2019), domain randomization has recently emerged as a particularly effective alternative Tobin et al. (2017). This approach systematically introduces variability into training environments—through perturbations in state parameters or environmental dynamics—encouraging agents to learn robust policies that remain effective under such variations.
This technique is particularly valuable for sim-to-real (sim2real) transfer, where it helps RL policies maintain performance when deployed in real-world environments, as demonstrated in robotics Zhao et al. (2020). In complex and dynamic domains such as RANs, where live system training is often impractical or prohibitively costly, domain randomization facilitates robust sim2real transfer. A complementary approach is sim-to-sim (sim2sim) skill transfer Gordon et al. (2019), which enables RL agents to transfer policies across simulated environments with varying characteristics, thereby enhancing generalization and increasing the likelihood of successful real-world deployment without repeated retraining Wang et al. (2021).
[…]
To address this challenge, this paper investigates generalization in RL for RAN applications, with a focus on improving adaptability and robustness under heterogeneous network conditions. The key contributions of this work are as follows:
  • We define RL generalization in the context of RANs, emphasizing the distinctive challenges compared to classical RL benchmarks.
  • We identify three enablers of generalization in RL for RAN: (i) accurate reconstruction of state representations that capture the behavior of RAN functions; (ii) increased diversity in training environments; and (iii) scalable distributed learning methods applicable to both simulated and operational networks.
  • We introduce a distributed learning architecture designed to address the dual challenge of ensuring data diversity and managing the computational complexity of training generalizable models. We show how this architecture supports large-scale data collection across simulated environments (for sim2sim transfer) and how its principles can extend to hierarchical, decentralized real-world RANs.
  • We develop a novel RL algorithm for link adaptation that demonstrates the impact of the proposed enablers on improving generalization.
  • We conduct extensive evaluations using a high-fidelity, system-level simulator compliant with 5G systems. Generalization is assessed via sim2sim transfer across environments with varying characteristics. Results show that the learned policies generalize to previously unseen radio conditions and outperform models fine-tuned to specific scenarios.
The paper is organized as follows. Sections 2 and 3 discuss the challenges and key enablers of RL generalization in RAN applications. Section 4 introduces graph models to enrich state representations with node attributes and topological information, while Section 5 describes a distributed learning architecture designed to enhance RL generalization in both simulated and live RANs. Section 6 details the link adaptation case study, and Section 7 outlines the corresponding RL design. Finally, Section 8 presents numerical evaluations, and Section 9 concludes the paper.
2 RL GENERALIZATION IN RAN APPLICATIONS
2.1 Definitions and Motivations
Following LeCun et al. (2015), we define RL generalization in the RAN context as the ability of a policy to perform well under new, previously unseen network conditions, deployments, and radio environments. We focus in particular on zero-shot generalization (ZSG) Kirk et al. (2023), which aims to produce RL policies that perform well in unseen RAN environments without additional training. This capability is critical because retraining for every environmental change (e.g., traffic pattern or network load variation) or deployment difference (e.g., different cells) is costly and impractical.
Complementary techniques such as few-shot adaptation and transfer learning can further facilitate RL deployment in RAN by enabling rapid adaptation from limited data or by leveraging knowledge from related scenarios. These methods are especially valuable in live systems, where RAN vendors—who develop and train the models—may lack direct access to field data or large-scale deployments, limiting their ability to evaluate or retrain policies under realistic conditions.
Achieving ZSG is also essential for scalable deployment of AI models in RAN systems. It enables network-wide replacement of a given function with a single, generalized policy, ensuring consistent performance and reducing the need for continual retraining. In contrast, relying on cell-specific fine-tuned models would degrade robustness and severely hinder scalability due to the increased complexity of lifecycle and data management. For example, maintaining separate models for thousands of radio cells for each AI-driven RAN function would be operationally burdensome and risk inconsistent behavior across the network.
2.2 Challenges
Key challenges in designing RL agents that generalize well in RANs include:
  1. 1.
    Heterogeneity of RAN environments: RAN deployments vary significantly in topology, scale, and radio access technologies. Differences in cell types (e.g., macro vs. small cells), hardware and deployment configurations (e.g., site location, antenna type and geometry), radio technologies (e.g., MIMO, mMIMO), and user distributions contribute to a highly diverse operational landscape. Such variability poses major obstacles to RL generalization in RANs.
  2. 2.
    Stochasticity and partial observability: Wireless channels are highly stochastic due to fading, interference, and random user activity. RL agents often operate under partial observability, relying on local measurements without full network visibility. This complicates credit assignment and reduces the reproducibility of experience—particularly in multi-cell settings, where actions may have delayed or indirect effects. Learning robust policies under uncertainty remains a significant challenge.
  3. 3.
    Non-stationary network dynamics: Wireless environments are inherently non-stationary. Factors such as user mobility, varying traffic loads, and time- and frequency-dependent channel conditions continually alter the environment’s transition and reward dynamics. Consequently, policies trained under specific traffic or mobility scenarios may degrade when deployed under sudden congestion, sparse demand, or changing user behavior. This temporal variability challenges static RL models and necessitates adaptive policy mechanisms.
Therefore, the RAN environment poses significant challenges to RL generalization, including highly variable operating conditions across time, space, and system configurations; partial and noisy observations; and frequent mismatches between training and deployment scenarios.
3 ENABLERS OF RL GENERALIZATION IN RAN
To promote RL generalization for RAN control functions, we identify three function-agnostic enablers: (i) robust state reconstruction (to handle partial and noisy observations), (ii) training-environment diversity (to cover a broad range of deployment scenarios), and (iii) distributed learning (to scale up data collection across many environments). In subsequent sections, we demonstrate their application to RL-based link adaptation.
3.1 Robust State Reconstruction
A prerequisite to attain model generalization over the RAN environment is robustness to uncertainty, that is, the model ability to cope with epistemic uncertainty arising from incomplete knowledge or limited data regarding the training domain Igl et al. (2019). Epistemic uncertainty, for example, occurs when certain areas of the input space are poorly represented or completely missing in the training data, preventing the model from making confident predictions. To ensure robustness under uncertainty for RL applications in RAN, the input state of a model must be carefully constructed to accurately reflect environmental changes, even when the available observations are limited, delayed, or noisy.
RAN functionalities typically operate under partial observability and heterogeneous information aging. Access nodes, for example, can only rely on local or incomplete views of the network state – either from direct sensing or from measurements reported by users or neighboring nodes. Furthermore, such information is inherently affected by uncertainties in its availability when a model needs it. For instance, while a model design may require certain measurements, such data may be temporarily unavailable, outdated, or unsupported depending on the capability of specific network nodes or user devices. Moreover, different observations needed to reconstruct the state of a RAN functionality may be gathered with very diverse timescales. As a result, the state reconstructed at a given time often contains features with different aging rates, leading to incomplete representations of the true state of the underlying system.
To enrich state reconstruction, we categorize information into static (e.g., deployment topology, configuration parameters), semi-static (or slowly varying, e.g., traffic load), and dynamic (or rapidly changing, e.g., sub-second radio channel conditions). Dynamic information pertains to measurable state dynamics specific to the RAN functionality that RL replaces. For example, for the link adaptation study case considered later, dynamic information includes path loss, channel state, and hybrid automatic repeat request (HARQ) feedback, and more. In contrast, static and semi-static information play a critical role in characterizing the radio environment, capturing the diversity of RAN deployments and enabling models to operate effectively across the network. Such information may describe network topology, site locations, orientations, inter-site relationships, transmission technologies (e.g., antenna array type, carrier frequency, bandwidth, transmit power), and user device characteristics (e.g., chipset type, capabilities), providing the context needed for models to handle a wide variety of devices.
This classification allows us to consider different approaches to reconstruct the state of a RAN function. The classic approach is to feed all information, indistinctly, to a deep Q-network (DQN) model. However, to improve generalization across diverse RAN deployments and propagation environments, static and semi-static information can be processed using graph-based methods, such as graph neural networkss Scarselli et al. (2009). Graph structures allow us to embed both radio access nodes and cell attributes, as well as their topological relationships, into rich state representations that can be transferred across different network deployments, improving model robustness to variations in network layout and environmental conditions Suárez-Varela et al. (2023); Shen et al. (2023); Lu et al. (2025). Section 4 further expands on this aspect.
3.2 Training-Environment Diversification
An effective approach to achieving ZSG in RL is domain randomization Zhao et al. (2020), which deliberately varies aspects of the training environment so that agents learn robust, transferable policies. By exposing the agent to conditions often beyond those expected in deployment—through randomized user locations, traffic loads, channel conditions, and interference levels—the learned policy overfits less and generalizes better. In this way, domain randomization narrows the sim2real gap by making simulations sufficiently diverse to capture real-world uncertainty.
In the context of RANs, domain randomization entails sampling heterogeneous network topologies, traffic and load scenarios, mobility patterns, and user densities. For example, an RL–based scheduler could train on randomized combinations of traffic demands and cell loads to improve generalization to unpredictable deployment conditions. Prior work in robotics has demonstrated the benefits of zero-shot transfer Zhao et al. (2020), and we argue similar benefits for RL in wireless networks, particularly when combined with domain adaptation techniques.
Despite its power, domain randomization poses three critical challenges:
  1. 1.
    Curse of dimensionality: As more parameters (e.g., user positions, mobility models, fading profiles, antenna patterns) are randomized, the parameter space expands exponentially. Sampling this space adequately becomes computationally expensive or infeasible.
  2. 2.
    Unrealistic extremes: Unconstrained randomization can yield non-physical scenarios—such as channel models with zero path loss or deployments involving hundreds of user equipments in a 100 meters cell—that misguide learning. To prevent this, domain knowledge and expert-defined constraints should be applied to keep randomization within physically plausible limits.
  3. 3.
    Training instability: Domain randomization can introduce high variance in reward signals, destabilizing training and leading to non-convergent policies, excessive exploration, or catastrophic forgetting. This instability can be mitigated through curriculum learning Graves et al. (2017); Matiisen et al. (2020), which gradually increases task complexity or environment randomness to ensure stable and progressive training.
3.3 Distributed Learning
Distributed learning improves model robustness in sim2sim transfer by enabling parallel exploration of diverse environmental configurations. Multiple actors interact with distinct randomized environments concurrently, achieving broader domain coverage and reducing wall-clock time compared to sequential sampling. Architectures such as IMPALA Espeholt et al. (2018), A3C Mnih et al. (2016b), Ape-X Horgan et al. (2018), and distributed PPO OpenAI et al. (2019) asynchronously update a shared policy from heterogeneous experiences, stabilizing training and mitigating overfitting to narrowly sampled conditions. A centralized replay buffer or gradient aggregator further integrates experiences across configurations, exposing the learner to a wider range of gradients and fostering robust, generalizable policies.
In Section 5, we present an architecture that scales distributed learning in RAN simulators to exploit environment diversification. We also discuss how its design principles can apply to live RAN deployments, where data diversity arises from the collective experience of distributed network nodes performing policy evaluation during training.
4 GRAPHS
[UE-centric network view.]Refer to caption [UE-centric graph.] Refer to caption
Figure 1: UE-centric representation of multi-cell relationships in cellular networks. (a) Radio environment snapshot as observed by a UE, reporting reference signal measurements from its serving cell (green) and its eight strongest neighboring cells (blue), each characterized by antenna configuration, bandwidth, and frequency. (b) Abstracted graph model in which the UE’s serving cell (central node, $h_{0}$) connects to neighboring cell nodes $h_{1}-h_{8}$. Edge thickness and node ordering represent measurement strength (e.g., RSRP), capturing inter-cell interference and handover affinities relevant to downstream machine learning tasks.
[…]
GNN extend traditional neural networks to graph-structured data, allowing the incorporation of connectivity information into the learning process. To achieve this, GNNs typically employ a feedforward architecture with multiple layers that iteratively aggregate and transform information characterizing the local neighborhood of a node using learnable permutation-invariant functions Scarselli et al. (2009). At each layer of a GNN, a node updates its embedding by combining its own features with a summary of its neighbors’ features, typically through operations such as message passing and neighborhood pooling Scarselli et al. (2009). The number of layers determines how far information propagates in the graph: for instance, one layer captures immediate neighbors, two layers include two-hop neighbors, and so on.
Several variants of GNNs have been proposed to capture different aspects of graph data, including:
[…]
    GCN: Extend convolutional networks to graph structures, effectively capturing local node interactions Kipf and Welling (2017).
[…]
    GAT: Incorporate attention mechanisms to weigh the importance of different neighbors during message passing Veličković et al. (2018).
  • Graph transformers: Leverage the self-attention mechanism of transformers to model long-range dependencies in graph-structured data, potentially capturing complex hierarchical relationships in RANs systems Ying et al. (2021).
Graph representations, such as those learned by GNNs, are particularly well suited for modeling RAN deployments, as they can capture both fine-grained local interactions (e.g., interference between nearby cells) and high-level global patterns (e.g., backbone connectivity or routing hierarchies) Lu et al. (2025). We adopt this approach to learn expressive network embeddings from static and semi-static network information that captures both site/cell attributes and inter-site/cell relationships within a RAN topology. Since GNN-based models share aggregation functions across all nodes and edges, the same model can be reused—without retraining—for graphs with varying topologies that represent spatial and relational patterns of different portions of a RAN deployment. Therefore, graph models facilitate better RL generalization in RAN by improving the robustness and transferability of the model across diverse deployment scenarios. Furthermore, because the information characterizing a RAN deployment is common across different functionalities, a learned embedding of the deployment can be reused in a variety of applications, including interference management, mobility management, and traffic prediction Shen et al. (2023).
Different graph representations of a RAN deployment can be constructed by modeling base stations, antennas, or user devices as nodes, while edges can capture communication links, interference relationships and geographical proximity. The graph structure in GNN-based models is often tailored to the specific use case. For example, Figure 1 illustrates a graph representing a UE-centric view of a RAN neighborhood, where the central node represents the serving cell of a UE, and surrounding nodes represent relevant interfering cells. Each node $i$ is associated with a feature vector $\mathbf{h}_{i}$ that characterizes the corresponding radio cell, while edges define proximity relationships between the associated cells.
The approach shown in Figure 1 is useful for designing RAN functionalities that operate on UE-specific parameters, such as link power control, transmission beamforming, or link adaptation, where the functional state varies per UE. In this case, the graph construction ensures that the context information provided to the GNN reflects the portion of the network topology most relevant to each UE—e.g., capturing relationships between surrounding cells that affect interference patterns, signal strength, and traffic conditions. In such a scenario, the serving cell can derive the graph instance using UE-side measurements (e.g., Reference Signal Receive Power (RSRP) values of neighboring cells), while additional context information can be exchanged via standardized interfaces, such as the Xn interface in 5G NR networks.
Section 8 describes how we integrate this approach with a DQN design for study case presented in Section 6.
5 DISTRIBUTED LEARNING
Since the introduction of the seminal DQN algorithm Mnih et al. (2013), several algorithmic and architectural advancements—such as GORILA Nair et al. (2015), A3C Mnih et al. (2016a), IMPALA Espeholt et al. (2018), Ape-X Horgan et al. (2018), R2D2 Kapturowski et al. (2018), Seed RL Espeholt et al. (2020), NGU Badia et al. (2020b), and Agent57 Badia et al. (2020a)—have progressively improved the scalability and effectiveness of distributed RL algorithms. Notably, IMPALA demonstrated how off-policy learning, combined with distributed actor-learner architectures, can effectively mitigate latency between action generation and gradient updates. These methods employ asynchronous updates, with distributed actors concurrently generating and evaluating training data across multiple parallel environments. This approach leads to more frequent policy updates, thereby accelerating learning and significantly improving performance. For example, the R2D2 architecture achieved up to a 20-fold increase in performance and a 50-fold reduction in training time compared to traditional DQN implementations.
5.1 Scaling Training in Network Simulators
In this work, we implement a large-scale distributed RL algorithm, illustrated in Figure 2. The algorithm is based on the distributed prioritized experience replay architecture proposed in Horgan et al. (2018), which decouples inference from learning, enables parallel inference, and significantly improves data collection capabilities. The architecture comprises three key components: a single learner operating on a graphics processing unit (GPU); multiple independent actors running concurrently on central processing units; and a centralized, yet sharded, prioritized replay memory. A sharded replay memory partitions the experience replay buffer into several smaller segments, or shards, each storing and managing a portion of the collected experiences independently—often distributed across multiple threads, processes, or machines—to enhance scalability and throughput. Each actor concurrently interacts with a group of simulations hosted by different CPUs, further accelerating the data collection process. Furthermore, each actor employs a distinct exploration policy to enrich the diversity of data. Experiences generated by each actor through interactions with its environment instances are asynchronously contributed to the replay memory shard.
The architecture is deployed on an high-performance computing (HPC) cluster, where the learner, actors, and replay memory are co-located on a single compute node, while groups of simulation environments execute in parallel across multiple additional nodes. This setup enables high-throughput distributed RL training with large-scale domain randomization over heterogeneous network simulations.
Figure 2: Overview of a scalable architecture for a distributed RL system integrated with wireless network simulators. A GPU-based learner updates neural network weights using batches sampled from a distributed replay memory containing experience trajectories. Updated weights are periodically shared with a set of CPU-based actors, which concurrently interact with diverse simulation environments, emulating variable channel conditions and user behaviors. The resulting experience streams are pushed to the replay memory, with sampling priorities updated based on the learner feedback.
5.2 Working Principles
The interaction among actors, shards, and the learner—illustrated in Figure 2—proceeds as follows:
Experience generation by actors: Each actor runs a replica of the current policy network on its assigned CPU core and interacts with multiple simulated environments, which are distributed across CPU cores on different compute nodes. At each step, or after completing an episode rollout, the actor generates experience tuples of the form (state, action, reward, next state). These experiences are first stored in a local buffer and then sent in batches to the replay memory, along with initial priority values.
Experience distribution to shards: Incoming experiences are assigned to replay shards using one of the following strategies:
  • Round-robin or load balancing: Experiences are distributed in a round-robin fashion or via a load-balancing mechanism to evenly fill all shards. For example, with four shards, an actor may send its first experience to shard 1, the next to shard 2, and so on cyclically. This strategy prevents bottlenecks and promotes uniform growth across shards.
  • Fixed actor-to-shard mapping: Alternatively, each actor—or group of actors—can be statically mapped to a specific shard, sending all experiences exclusively to that shard throughout training.
Data storage in shards: Each shard operates as an independent, typically circular, replay buffer. When full, a shard evicts older experiences or, in prioritized replay, replaces those with the lowest priority. Since shards are decoupled, data insertions to one shard do not interfere with others, enabling parallel writes.
Sampling from shards by the learner: The probability of sampling a transition $e_{i,j}$ from the $j^{\text{th}}$ shard $\mathcal{R}_{j}$ is:
$\Pr(e_{i,j})=\frac{p_{i,j}^{\alpha}}{\sum_{j}\sum_{k}p_{k,j}^{\alpha}}=\left(\frac{p_{i,j}^{\alpha}}{\sum_{k}p_{k,j}^{\alpha}}\right)\cdot\left(\frac{\sum_{k}p_{k,j}^{\alpha}}{\sum_{j}\sum_{k}p_{k,j}^{\alpha}}\right)=\Pr(e_{i,j}\mid\mathcal{R}_{j})\cdot\Pr(\mathcal{R}_{j}).$ (1)
The learner periodically queries each shard to retrieve its total priority and compute $\Pr(\mathcal{R}_{j})$, as defined in Equation 1. It then samples a proportional number of experiences from each shard based on $\Pr(e_{i,j}\mid\mathcal{R}_{j})$, combines them, and shuffles to form a training batch. This strategy ensures batch diversity by drawing across all shards.
Learning update: The learner performs a gradient update on the neural network using the sampled batch. After each update, the learner independently updates the priorities of the sampled experiences.
Priority updates: Each sampled experience is associated with a priority, typically derived from the temporal difference (TD) error. After updating priorities, the learner sends these values to the corresponding shard. Each shard independently maintains its own data structure (e.g., segment tree or heap) to manage priorities and facilitate efficient sampling.
Iterative process: Actors continuously generate new experiences and send them to their designated shards, while the learner repeatedly samples from shards and updates the policy. This pipeline runs in parallel. To maintain policy consistency, the learner periodically broadcasts updated network weights to all actors. As a result, the replay memory contains a mixture of experiences collected under different policies, making the learning process inherently off-policy.
5.3 Integrating distributed learning in RAN systems
Figure 3: An AI architecture that integrates distributed learning in a 5G system architecture, with a learning engine realized as either a standardized or a standalone support function.
The foundational design principles of RL architectures—distributing policy evaluation (i.e., data generation) across actors while centralizing training and data management—align naturally with the hierarchical structure of RAN systems. These principles support an AI architecture in which a centralized learning engine orchestrates and provides support services to AI applications deployed in distributed RAN network functions, such as gNB centralized units, gNB distributed units, and x/rApps, which are responsible for various control and prediction tasks, as illustrated in Figure 3.
The architecture in Figure 3 enables distributed learning by allowing RAN NFs that host AI-driven functionalities to collect experiences from local environments and contribute to a shared dataset for training a generalized model. Analogous to domain randomization in simulations, distributing policy evaluation across RAN NFs during training increases sample diversity, mitigates data bias, and enhances robustness against uncertainties inherent to real-world RANs.
This architecture features a learning engine deployment that is agnostic to specific AI applications, while the deployment of actors (or inference functions more broadly) is use-case dependent. For example, actors replacing physical layer (L1) and medium access control layer (L2) functions—such as resource scheduling or link adaptation—could be hosted in gNB-DUs. In contrast, actors replacing networking layer (L3) functions—such as load balancing or mobility handover—could reside in gNB-CUs. To support AI-native network orchestration functions—e.g., optimizing the coverage-capacity trade-off across a network area—actors may be realized as rApps. Alternatively, actors (or inference functions) may be deployed as xApps in the nea-real-time RAN intelligent controller (RT RIC) of an ORAN system, providing insights or configuration updates to RAN NFs.
To fulfill its central role and remain use-case agnostic, the learning engine can be deployed at higher layers of the RAN architecture. A cloud-based implementation provides abundant and flexible compute and storage resources to deliver AI services to distributed RAN functions. The learning engine supports a variety of services, including model training, data services (e.g., data generation control, storage, and management), and lifecycle management (LCM) operations that ensure model compliance with reliability, performance, security, ethical, and standards requirements. While LCM is centrally orchestrated, certain computational tasks—such as performance monitoring or drift detection—can be delegated to RAN NFs that execute AI models locally. This approach concentrates computationally intensive tasks, such as training and large-scale data storage, where they are most needed, thereby promoting a cost-efficient and scalable integration of AI into RAN systems.
We consider two approaches to integrating a learning engine into RAN: (a) as a standalone support function, and (b) as a standardized function.
Figure 3 exemplifies a standalone learning engine coexisting with an service management and orchestration (SMO) platform in the network automation layer, communicating with actors (inference functions) hosted in RAN NFs via proprietary or extended standardized interfaces. The ML-D interface (shown in green) supports data collection from distributed RAN NFs; ML-Cd (shown in black) orchestrates training data generation for specific AI applications; and ML-Cm (shown in blue) supports artifact management and model deployment. A standalone approach offers maximum design flexibility, allowing RAN NFs and r/xApps to locally process raw data into customized training samples that meet algorithmic requirements—such as state-action-reward tuples for RL—prior to transferring them to the learning engine.
A standardized approach, on the other hand, relies on standardized interfaces and functional frameworks. For example, using ORAN terminology, Figure 3 illustrates a learning engine implemented as an SMO function (SMOF), relying on other SMOFs, such as a data management entity (DME), for data collection, management, and storage. This approach mandates the use of standardized protocols and formats, which may not always align with specific AI training requirements.
For instance, RAN NFs may transmit raw network observations to a DME over O1/O2 interfaces (shown in red) in accordance with interface specifications. However, these raw observations may not be directly usable for training. The DME must then process the data into training samples suited to the AI application, which are subsequently transmitted—e.g., via the R1 interface (shown in green)—to the learning engine. The primary motivation for a standardized approach is to enable inter-vendor interoperability, allowing network operators to build architectures composed of components from different vendors. However, standardizing the learning engine’s functionalities may impose significant design constraints.
6 CASE STUDY: LINK ADAPTATION
[…]
LA is a core function of modern wireless communication systems that employs adaptive modulation and coding to maximize spectral efficiency. Specifically, link adaptation (LA) optimizes the modulation order and code rate of a packet transmission to match the link capacity based on the underlying radio conditions, using receiver-side channel state information (CSI) such as channel quality indicator (CQI3rd Generation Partnership Project (3GPP). The LA parameters are then encoded into a unique value, known as the modulation and coding scheme (MCS) index, and sent to the receiver to facilitate packet decoding 3rd Generation Partnership Project (3GPP).
6.1 Outer-Loop Link Adaptation
[…]
Although several variants of this approach have been proposed Buenestado et al. (2014); Durán et al. (2015); Park et al. (2015); Blanquez-Casado et al. (2016); Ohseki and Suegara (2016); Delgado et al. (2017); Nemeth et al. (2021); Ramireddy et al. (2022); Nicolini et al. (2023), factors such as user mobility, traffic burstiness, channel aging, limited receiver-side information, and heterogeneous UE populations (e.g., with varying receiver and hardware capabilities) make LA a difficult control task to model Chen et al. (2023), rendering OLLA suboptimal.
6.2 Reinforcement Learning-Based Link Adaptation
Recognizing the limitations of traditional heuristics, there is growing interest in developing alternative, proactive solutions to LA by leveraging the predictive capabilities of machine learning (ML), and RL in particular. Numerous studies have proposed LA solutions for 5G NR networks using various Q-learning variants. Notably, Leite et al. Leite et al. (2012), Bruno et al. Bruno et al. (2014), and Mota et al. Mota et al. (2019) employed tabular Q-learning; Wu et al. Wu et al. (2020), Chen et al. Chen et al. (2023), and Kela et al. Kela et al. (2022) utilized DQN; while Geiser et al. Geiser et al. (2022) proposed an actor-critic RL algorithm for optimizing MCS selection. Other work, such as Saxena et al. Saxena et al. (2021) and Karmakar et al. Karmakar et al. (2017), has explored bandit-based approaches.
While these studies demonstrate that RL can improve LA performance over rule-based methods, they introduce new challenges that limit practical viability. For example, tabular Q-learning does not scale well with the state-action space of LA, and bandit-based approaches often rely on overly simplified problem formulations that ignore retransmissions. Most notably, many studies train and evaluate RL models using simplistic simulators that assume ideal conditions and fixed environments, without considering whether the models can generalize effectively to the diverse conditions found across radio cells in a real-world, network-wide RAN deployment.
7 MODELLING LINK ADAPTATION AS AN MDP
We jointly model the downlink LA and HARQ processes in RANs as a MDP, and present a novel DQN-based design combined with the enablers described in Section 3 to demonstrate RL generalization for RAN applications.
7.1 Episode design
Link adaptation and HARQ operate on a per-UE and per-packet transmission basis. We formulate this problem as an episodic MDP Sutton and Barto (2018) defined by $\mathcal{M}=\left\langle\mathcal{S},\mathcal{A},p,r,\gamma,\rho_{0}\right\rangle$, where $\mathcal{S}$ denotes the state space, $\mathcal{A}$ the action space, $p(s^{\prime}\mid s,a)$ the transition dynamics, $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ the reward function, $\gamma\in[0,1)$ the discount factor, and $\rho_{0}$ the initial state distribution.
An episode models the lifespan of a UE packet in the HARQ process—from its first transmission to either a successful reception or the packet being dropped upon $N$ transmission attempts, as illustrated in Figure 4. This enables us to train a single RL policy from the collective experience generated by any UEs across the network. Each step in the episode represents the duration of a packet transmission in the HARQ process, from the selection of LA parameters (i.e., the RL action) to the reception of the associated HARQ feedback, i.e., a positive acknowledgment (ACK) or a negative acknowledgment (NACK) for successful or failed transmission, respectively. For instance, the 3GPP 5G NR system, used in our evaluations, supports at most four packet retransmissions. Hence, the episode length $N$ may range from one to five steps. Each step is characterized by a state, an action, and an associated reward, as presented next.
Figure 4: Markov decision process formulation for downlink link adaptation and HARQ in wireless communication. The figure illustrates how episodes in a MDP are structured to model the interaction between the BS and UE during the downlink transmission and hybrid HARQ protocol. Each episode corresponds to the transmission of a single packet and is composed of a sequence of steps, each defined by a state ($s_{t}$), an action ($a_{t}$), and a reward ($r_{t}$). The process begins with the arrival of a new packet at the BS and proceeds through successive transmissions, receiving NACKs or ACKs from the UE. Colored segments represent repeated transmissions of the same packet due to decoding failures, until successful reception is confirmed via an ACK. This MDP-based abstraction captures the sequential decision-making dynamics inherent in adaptive modulation and retransmission strategies.
As the goal of LA is to adapt the MCS parameters to time- and frequency-varying channel conditions to maintain high link spectral efficiency, we propose a reward design that reflects the spectral efficiency of individual transmissions, as
$r_{n}=SE_{n}\cdot\delta_{n}-\alpha\cdot n\cdot(1-\delta_{n})\qquad n=1,\dots,N,$ (2)
and
$\delta_{n}=\begin{cases}1&\text{if ACK at $n$-th transmission attempt},\\ 0&\text{if NACK at $n$-th transmission attempt},\end{cases}$ (3)
where $n$ denotes the packet transmission counter, the $SE_{n}$ indicates the spectral efficiency of the packet upon transmission attempt $n$, while $\alpha\geq 0$ is a scalar robustness weight used to penalize packet retransmissions.
Optimizing exclusively for spectral efficiency can result in a large number of retransmissions and, thus, lower throughput (which accounts for the time required to deliver a packet). Rule-based LA algorithms like OLLA (cf. Section 6) tackle this by requiring one to adhere to a specific BLER target. Conversely, the reward design in Equation 2 integrates two mechanisms that allow learning policies that discourage exploiting retransmissions to improve spectral efficiency at the expense of throughput. Firstly, upon a successful packet transmission (ACK) at the $n$-th attempt, the value of $SE_{n}$ accounts for the cumulative time-frequency resources used to deliver the data payload with all $n$ transmission attempts. Furthermore, the reward for each failed packet transmission (NACK) is a penalty that increases with each transmission $n$ scaled by $\alpha$.
The reward design Equation 2 is agnostic to the underlying radio transmission technology, scheduling policy, and system bandwidth—a key feature to pursue model generalization over the RAN environment that allow us to deploy a single model in heterogeneous RAN systems with network nodes operating with bandwidth varying from a few to hundreds of MHz. Specifically, using an expression of spectral efficiency renders the reward agnostic the transport block size (TBS) (i.e., information bits) carried by the packet and the number of resource elements $N_{RE}$ allocated by the scheduling policy. Unlike the range of values of TBS and $N_{RE}$, the nominal link spectral efficiency values is agnostic to the system bandwidth and regulated by communication standards such as in the 3GPP NR MCS tables 3rd Generation Partnership Project (3GPP). Furthermore, considering the spectral efficiency per spatial layer further allows us to train a single model that can serve both single-layer and multi-layer MIMO transmissions.
7.2 Action space design
At its essence, the LA action space consists of a set of MCS index values $\mathcal{A}=\{0,\dots,M-1\}$ reflecting a set of discrete data rates supported by a communication standard. In our evaluations, for instance, we consider the MCS index values defined in the 3GPP 5G NR specifications (e.g., Table 5.1.3.1-1 or Table 5.1.3.1-2 in 3rd Generation Partnership Project (3GPP), supporting modulations orders up to 64QAM and 256QAM, respectively). An action $a_{m}=m$ implicitly provides the modulation order, code rate, and link spectral efficiency to be used for a packet transmission, given a state $s\in\mathcal{S}$. This information, combined with the allocation of time-frequency resource blocks to the UE, enables to determine the TBS used for a new packet transmission according to NR specifications 3rd Generation Partnership Project (3GPP, 3GPP), which can then be used to evaluate the corresponding reward according to Equation 2. Packet re-transmissions, however, reuse the same TBS of the initial packet transmission, while the modulation order and code rate can change.
7.3 State space design
To enable distributed learning of a single RL policy from the collective experience of all UEs across the network, we model the state of the LA process by combining (a) a UE-centric view of the RAN environment surrounding the UE consisting of semi-static information characterizing network nodes, radio cells and the UE itself (cf. Section 4), and (b) dynamic information characterizing relevant features of the LA process. Together with domain randomization during training (cf. Section 3), this allows us to training models that generalize over the RAN environment.
7.3.1 Semi-static information
We describe the network deployment surrounding a UE considering its serving cell and subset of most relevant interfering cells, as illustrated in Figure 1, consisting of $K_{I}$ cells whose signal strength measured by the UE exceeds a threshold. As such, the cells comprising the UE’s surrounding may change over time depending on the UE mobility. For each cell, we consider semi-static configuration information, such as the site type (e.g., LTE, NR, etc.), carrier frequency, antenna array configuration (e.g., MIMO or mMIMO), geometry information such as inter-site distance, bandwidth and transmit power configurations, etc.
Furthermore, as the strength and the quality of the signals perceived by a UE can be significantly affected by manufacturing characteristics, such as chipset and antennas hardware and software quality (e.g., algorithms for signal processing and channel state estimation), the MDP state needs to capture information characterizing the UE itself to achieve generalization with respect to the diversity of the UE population. To this end, we consider an indicator representing the UE type and size of the UE antenna array.
In our design, we feed this information either directly to a DQN network or graph-based model (as introduced in Section 4). While a limited set of semi-static parameters suffice to characterize the RAN deployment in our numerical evaluations (cf. Section 8), additional information may be beneficial for practical deployment in live networks.
7.3.2 Dynamic information
Furthermore, we consider a set of observable information relevant to the LA process of a UE prior to a packet transmission that evolve with fast dynamics and include measurements of the radio link condition, resource availability, and UE state. As LA operates on a sub-millisecond timescale, information aging plays a crucial role to confidently predict the best MCS index to apply for a packet transmission. We therefore apply either a soft- or a hard-aging model to different types of information, as described next.
Packet transmission counter: As LA operates on a per-packet basis, and packet retransmissions may exploit previously received copies of the packet to improve the decoding probability, we include the packet transmission counter $n$ (cf. Equation 2) as part of the state features. The ability to distinguish the first transmission and individual retransmissions helps to learn policies that can optimize LA parameters across multiple transmissions of the same packet.
UE-specific measurements: The 3GPP NR system supports a variety of signals and measurements to observe channel conditions for different purposes. Here, we consider exclusively observations that can be available prior to a packet transmission, such as RSRP measurements and CSI reports. RSRP measurements provide a surrogate to pathloss, which reflects macro channel conditions (e.g., shadowing) and distance between the UE and its serving cell. CSI reports, consisting of a CQI value and a suggested transmission rank, represent the channel dynamics observed by the UE. While the reported CSI typically comply with technical specifications to allow inter-vendor interoperability, differences in UE hardware and proprietary algorithms used to derive CSI reports, can cause UEs of different manufactures to report different CSI observations for the same channel conditions.
Historical CSI: Rather than using a single CSI report, we consider a sequence of $K_{CSI}\geq 1$ historical CQIs and rank values, as well as a weighted average CQI, where the individual CQIs are weighted based on the associated rank.
To account for information aging, we consider a time window $W$ and override CQI and rank values to $-1$ for CSI reports older than $W$. The time window $W$ can be chosen, for instance, based on the channel coherence time, thus allowing to learn that too old information is less relevant to derive LA parameters for the current packet transmission.
UE buffer: The UE buffer status provides insights about data awaiting to be transmitted. Rather than using directly the UE buffer size $B_{t}$ at time $t$, we scale $B_{t}$ by the maximum transport block size $TBS_{\max}$ allowed by the system as
$b_{t}=\min\left\{N_{TTI},\;\frac{B_{t}}{TBS_{\max}}\right\},$ (4)
thereby expressing the buffer size in terms of number of transmissions that would be required to flash the UE buffer, assuming the most favorable channel and resource availability conditions, with $N_{TTI}$ denoting a suitable upper bound. This expression renders information about UE buffer state agnostic to the system bandwidth and supported modulation orders, thus enabling re-usability in heterogeneous RANs.
Expected spectral efficiency (SE): We further consider as part of the LA state an estimate of the link spectral efficiency derived from the most recent historical CSI reports. Specifically, for a given UE $i$, we maintain an average spectral efficiency defined as
$SE_{i}=\frac{\sum_{k=1}^{K_{CSI}}SE_{k}^{(i)}\cdot rank_{k}^{(i)}}{\sum_{k=1}^{K_{CSI}}rank_{k}^{(i)}},$ (5)
where $SE_{k}^{(i)}$ and $rank_{k}^{(i)}$ denotes the link spectral efficiency and the transmission rank, respectively, associated to the $k$-th historical CSI report available for UE $i$. In particular, we model $SE_{k}^{(i)}$ as the nominal spectral efficiency value associated to the $k$-th CQI value reported by the UE as defined, for instance, by the 3GPP 5G NR specifications 3rd Generation Partnership Project (3GPP).
HARQ feedback: Link adaptation algorithms traditionally rely on UE HARQ feedback (i.e., ACK or NACK) to adjust internal link quality estimates for the next transmission. Such information remains relevant also to our design.
We consider a sequence of $K_{H}$ historical HARQ values $\{h_{k}^{(t_{k})}\}_{k=1}^{K_{H}}$, where the superscript $t_{k}$ denotes the time when the $k$-th HARQ value was reported. We model $h_{k}^{(t_{k})}$ as equal 1 if an ACK is reported at time $t_{k}$ and $h_{k}^{(t_{k})}=-1$ for NACK.
However, the relevance of a HARQ value $h_{k}^{(t_{k})}$ to reconstruct the MDP state for a transmission at time $t>t_{k}$ strongly decays with its age $t-t_{k}$. For example, for bursty or ultra-reliable low latency (URLLC) traffic, or in conditions of high cell load, two consecutive transmissions to the same UE may occur with a wide temporal gap (e.g., from tens or hundreds of milliseconds to seconds). As such, a HARQ value $h_{k}^{(t_{k})}$ whose age largely exceeds the channel coherence time may not help reconstruct the current state of the link.
Therefore, rather than using directly the HARQ value $h_{k}^{(t_{k})}$, we define a soft-aging model where $h_{k}^{(t_{k})}$ is linearly scaled based on is age $t-t_{k}$ and a maximum aging $W$ as
$\tilde{h}_{k}^{(t_{k})}=\begin{cases}h_{k}^{(t_{k})}\left(1-\frac{\max\{0,\;t-t_{k}\}}{W}\right)&\text{if}\;t-t_{k}\leq W\\ 0&\text{otherwise,}\end{cases}$ (6)
The soft-aged historical HARQ $\tilde{h}_{k}^{(t_{k})}$ takes values within the interval [1, 1]11[-1,\;1][ - 1 , 1 ], with $\tilde{h}_{k}^{(t_{k})}$ starting from -1 for NACK or 1 for ACK, respectively, and gradually reaching zero as the HARQ age approaches $W$. A value $\tilde{h}_{k}^{(t_{k})}=0$ indicates that the HARQ feedback is too old to provide reliable information for selecting LA parameters for the current transmission.
Historical LA parameters: Link adaptation parameters used in prior transmissions for a UE (e.g., previous MCS index and rank) can provide insights to reconstruct the MDP state for selecting the UE’s LA parameters at time $t$. We therefore consider a sequence of $K_{LA}$ historical MCS index values $\{m_{k}^{(t_{k})}\}_{k=1}^{K_{LA}}$ and historical of rank values $\{l_{k}^{(t_{k})}\}_{k=1}^{K_{LA}}$. To account for information aging we apply a hard-aging model where historical information older than the maximum age window $W$ is assigned a special value (namely -1) as
$\tilde{m}_{k}^{(t_{k})}=\begin{cases}m_{k}^{(t_{k})}&\text{if}\;t-t_{k}\leq W\\ -1&\text{otherwise.}\end{cases}$ (7)
and
$\tilde{l}_{k}^{(t_{k})}=\begin{cases}l_{k}^{(t_{k})}&\text{if}\;t-t_{k}\leq W\\ -1&\text{otherwise.}\end{cases}$ (8)
8 EXPERIMENTAL EVALUATIONS
8.1 Training Setup
8.1.1 Distributed Training Architecture
Our training system, illustrated in Figure 2, utilizes a distributed training architecture comprising a single learner, 40 actors, and a shared replay memory composed of several independent shards. These components are co-located on a shared compute node, while the simulation environments are hosted on separate compute nodes to facilitate distributed and scalable experience collection. To maximize throughput, the system operates asynchronously, allowing actors and the learner to work independently without global synchronization.
The replay memory is partitioned into four independently prioritized shards, each maintaining its own priority queue. Each shard has a capacity of 4 million transitions, enabling long-term experience storage and sampling diversity. Actors write transitions asynchronously to specific shards in batches, reducing communication overhead and supporting scalable data ingestion, similar to Kapturowski et al. (2022).
As shown in Figure 5, each actor steadily writes batches to its assigned shard at a rate of approximately 18–20 batches per minute after an initial ramp-up period. This time series, covering a representative 30-hour training run, displays data from four actors (IDs 0, 10, 20, and 30) and highlights both the system’s ability to reach a stable ingestion regime and the even distribution of load across actors and shards. The close alignment in write rates across disparate actor indices provides strong empirical evidence of the scalability and load-balancing capabilities of our distributed experience collection infrastructure.
Figure 5: Per-actor replay-memory ingestion rates during large-scale distributed training. Each curve shows the number of 500-sample batches written per minute by four representative actors (IDs 0, 10, 20, and 30) to their assigned shards over a 30-hour run. After an initial ramp-up and shard initialization jitter (14 May, 17:00–19:00 UTC), all actors converge to a steady rate of 18–20 batches/minute. The uniformity of these curves confirms consistent ingestion throughput and balanced load across the actor population.
Actors run local replicas of the policy network and interact concurrently with 14 environment instances. [^1] Section 10.2 provides additional computation and deployment details. Similar to Horgan et al. (2018), each actor $i\in\{0,\cdots,39\}$ executes an $\epsilon_{i}$-greedy policy, defined as $\epsilon_{i}=\epsilon^{1+\frac{i}{N-1}\alpha}$, where $\epsilon$ is set to 0.4 and $\alpha$ to 8.5. This exploration policy remains fixed throughout training. On average, each actor generates approximately 132 transitions per second. Incoming transitions are buffered locally before being dispatched asynchronously to replay memory in batches of 500 transitions. Prior to training, a warm-up phase collects 45,000 transitions—without gradient updates—to populate the replay memory with a diverse initial dataset.
The learner is responsible for computing gradients and updating model parameters. It prefetches up to 16 batches of 512 transitions each from the replay memory shards. This prefetching ensures a continuous supply of data for gradient computation. On average, the learner processes 100 batches per second, corresponding to approximately 51,200 transitions per second. Model parameters are updated and broadcast asynchronously to all actors every 512 learner steps (roughly every 5 seconds), allowing actors to operate with relatively fresh policies while maintaining a decoupled learning and acting process.
The data flow proceeds as follows: actors collect environment transitions, locally buffer them, and push them to replay memory shards. The learner samples from these shards, performs gradient updates, and propagates the updated parameters back to the actors.
Table 1: RAN environment simulation parameters. Bold values are default parameters used in testing scenarios (which randomize over seeds only).
Parameter Value range Description
Duplexing type TDD Fixed
Carrier frequency 3.5 GHz Fixed
Deployment type 1-site 3-sector
Site type {MIMO, mMIMO} Randomized
Antenna array 1x2x2 MIMO (4) Fixed
8x4x2 mMIMO (64) Fixed
Cell radius {166, 300, 600, 900, 1200} m Randomized
Bandwidth {20, 40, 50, 80, 100} MHz Randomized
Number of sub-bands {51, 106, 133, 217, 273} Randomized
DL TX power {20, 40, 50, 80, 100} W Randomized
UE antennas {2, 4} Randomized
Max transm. rank {2, 4} Randomized
Max num. transm. 5 Fixed
UE traffic type {FB, eMBB} Randomized
Number FB UEs {1, 5, 10} Randomized
Number MBB UEs {5, 10, 30, 50, 100} Randomized
Speed UE FB {0.67, 10, 15, 30} m/s Randomized
Speed UE eMBB {0.67, 1.5, 3} m/s Randomized
UE receiver types {type0, type1, type2, type3} Randomized
Indoor probability {0.2, 0.4, 0.8} Randomized
8.1.2 Training Environment Diversification
We train the RL policy to generalize in the RAN environment by exposing it to a wide range of deployment scenarios, radio conditions, traffic patterns, interference levels, and load distributions, as discussed in Section 3.2. While such diversity is naturally present in live RAN systems—where training data can be generated across network nodes—we use domain randomization to recreate comparable conditions within network simulators.
Table 2: Benchmark network configuration scenarios used for testing.
Name Type Description Traffic User Types
B1 Single-cell single UE 1-site 1-cell MIMO, 1 outdoor UE at different locations FB Homogeneous (type0)
B2 Stable interference 1-site 3-cell (a. MIMO, b. mMIMO), 80% indoor UEs FB Homogeneous (type0)
B3 Dynamic interference 1-site 3-cell (mMIMO), 80% indoor UEs eMBB Homogeneous (type0)
B4 High mobility 1-site 3-cell (mMIMO), 0% indoor UEs FB Homogeneous (type0)
B5 Realistic traffic 3-site 9-cell (mMIMO), 80% indoor UEs FB and eMBB Homogeneous (type0)
Each training scenario is generated using an event-driven, 5G-compliant system simulator that operates in time division duplexing (TDD) mode on a 3.5 GHz carrier frequency, with single-user multiple input multiple output (SU-MIMO) transmission and a physical layer model based on numerology $\mu=0$ as specified in 3GPP NR standard 38.211 3rd Generation Partnership Project (3GPP).
A training scenario consists of three trisectorial radio sites, each randomly configured as either MIMO or mMIMO. Site configurations are randomized with respect to cell radius, system bandwidth, and downlink transmit power, using parameters sampled from Table 1. Further randomization is applied to cell load, traffic type, indoor/outdoor UE distribution, and UE receiver types. A UE generation process randomly selects the number of UEs with full buffer (FB) and eMBB traffic to be deployed, drawing from a predefined set of indoor/outdoor probabilities listed in Table 1.
Each eMBB UE generates traffic based on a packet arrival process and a packet size distribution derived from historical field data, with payloads ranging from 20 to approximately 1000 Kbytes. Each UE configuration is then randomized by the number of antennas, mobility speed, and receiver type. The latter emulates variations in channel estimation accuracy associated with different UE chipsets under the same channel conditions.
For hyperparameter tuning, we train RL policies on 40,000 training scenarios generated from 160 random seeds and 250 randomized simulation parameter configurations, as defined in Table 1. Each scenario simulates five seconds (equivalent to 5,000 transmission time intervals), producing approximately 16,000 training samples on average and yielding a total of roughly 650 million samples across the full hyperparameter sweep.
For the general performance evaluations presented in Section 8.3, we employ a larger dataset of 240,000 training scenarios generated from the same 160 seeds but with 1,500 randomized parameter configurations, also drawn from Table 1. This setup produces approximately 3 billion training samples in total.
[Cell throughput relative gains across benchmark scenarios.] Refer to caption [Spectral efficiency relative gains across benchmark scenarios.]Refer to caption
Figure 6: Performance comparison of specialized and generalized RL-based LA policies across five benchmark scenarios (B1–B5). The boxplots illustrate relative gain or loss (%) over a state-of-the-art baseline for two RL-based LA policies, both employing the same neural network architectures—a six-layer multilayer perceptron with 256 neurons per layer. The policies differ only by training procedure: a specialized policy trained exclusively for full-buffer (FB) conditions (red), and a generalized policy trained across varied network conditions (blue). Results are presented for five benchmark scenarios (B1–B5), showing the distribution of relative performance gains in: (a) downlink cell throughput, and (b) spectral efficiency. Each boxplot displays the median (diamond), interquartile ranges (box), whiskers extending to 2×2$\times$2 ×IQR, and outliers (circles).
8.2 Test Setup
For testing purposes, we define five benchmark scenarios, described in Table 2, whose conditions are not directly encountered during training. Benchmark B1 consists of a single-cell, single-UE deployment positioned randomly and operating in MIMO transmission mode. This scenario represents a challenging outlier for evaluating the link-curve performance of RL policies against the state-of-the-art OLLA algorithm used in 5G networks, which is specifically optimized for such conditions.
Scenarios B2–B5 consist of homogeneous network deployments—either MIMO or mMIMO—with simulation parameters based on the default values indicated in bold in Table 1. Each benchmark targets specific testing conditions: B2 evaluates stable interference under full-buffer traffic using either MIMO (B2(a)) or mMIMO (B2(b)) deployments; B3 simulates dynamic interference in mMIMO deployments due to eMBB traffic; B4 assesses performance for UEs traveling at high speeds (100–200 km/h); and B5 evaluates performance under realistic traffic conditions involving a mix of FB file transfers and chat traffic types not used during training.
8.3 General Results
To validate our design concepts for RL generalization in RAN applications, we compare the performance of two LA policies trained using the same model architecture and hyperparameters but under different data regimes. The first policy is trained to generalize using randomized network scenarios, while the second is specialized for FB traffic conditions by using an equal amount of simulated networks based on benchmarks B2(a), B2(b), and B4, with randomized seeds. In both cases, we use an MLP with six hidden layers and 256 neurons per layer. Inputs are described in Section 7.3, and the output layer comprises 28 nodes, each representing the Q-value associated with an MCS index compliant with 5G specifications (Table 5.1.3.1-2 in 3GPP TS 38.214 3rd Generation Partnership Project (3GPP)). Both policies are trained using a robustness weight $\alpha=0.5$ (see Equation 2).
We compare both policies with a state-of-the-art outer-loop LA design that targets a $10\%$ BLER, representing the most common approach adopted in real-world 5G NR systems.
Figure 6 summarizes the test results obtained across the five benchmarks described in Table 2, reporting statistics of gains and losses in average cell throughput (see Figure 6) and spectral efficiency (see Figure 6) relative to the baseline.
First, we observe that the generalized RL policy significantly outperforms the 5G baseline across all benchmarks, achieving gains in average cell throughput and spectral efficiency close to or exceeding 10% in FB MIMO and mMIMO scenarios, and surpassing 20% in high-mobility conditions.
Second, under FB traffic conditions (i.e., benchmarks B2 and B4), the generalized policy performs on par with the specialized policy, while significantly outperforming it in B3 (eMBB traffic) and B5 (mixed traffic), with gains of approximately 4× and 2×, respectively. In benchmark B1—analyzed in more detail in the following subsection—the generalized model also achieves higher gains with lower variance in both throughput and spectral efficiency.
These results suggest that RL algorithms trained for generalization using the enablers proposed in this paper can outperform both heuristic baselines and models optimized for narrow, scenario-specific conditions. Unlike specialized models, which perform well only under conditions seen during training and often lack robustness, the generalized approach proposed here provides the adaptability needed for scalable RL deployment in realistic networks, reducing the need for frequent retraining.
Additional evaluations on hyperparameter tuning—including model architecture, layer normalization, and loss functions—are presented in Section 10.3.
[Average throughput gains in benchmark B1.] Refer to caption [Average MCS index selected in benchmark B1.]Refer to caption
Figure 7: Performance analysis in single-cell single-UE benchmark B1, with the model tested on several realization of the scenario with different UE positions ranging from 40 to 1250 meters from the serving cell. Throughput gains are measured with respect to the OLLA baseline.
8.3.1 Link-Level Performance
Figure 7 further examines the performance of both models in the single-cell single-UE benchmark B1. This benchmark is relevant for two main reasons: it reflects outlier conditions never encountered during training, and it represents the scenario for which the OLLA baseline is optimized. The latter provides reference values for link-level spectral efficiency used, for instance, by 3GPP to standardize CQI and MCS tables 3rd Generation Partnership Project (3GPP).
We evaluate both the generalized and specialized models against the baseline across several initial UE locations, ranging from 40 meters to 1250 meters from the serving base station. At each location, we conduct 50 simulations using distinct random seeds for channel and environment realizations. Figure 7 shows that the generalized RL model consistently outperforms both the baseline and the specialized model at all locations. While average throughput gains range between 2% and 5.5%, it is important to note that the OLLA baseline nearly reaches link capacity under B1 conditions, leaving limited room for improvement.
Figure 7 shows that the generalized RL model achieves these gains by selecting, on average, slightly more conservative MCS indices at greater distances from the base station. This conservatism reduces the number of retransmissions experienced by the RL policy, resulting in modest but consistent throughput gains. In contrast, the specialized model is overly conservative: near the base station, it fails to recognize ideal channel conditions (i.e., low path loss attenuation), instead selecting lower MCS indices. This results in fewer information bits being transmitted and thus a performance loss. At other distances, this model also tends to select slightly more conservative MCS indices, yielding smaller gains over the baseline.
8.4 Reward Analysis
Figures 8 and 9, together with the reward definition in Equation 2, highlight the dual role of the robustness weight $\alpha$ in the design of RL-based link adaptation.
Specifically, we trained three generalized policies using the same model architecture, hyperparameters, and training procedure as in the previous section, but with different settings of $\alpha$: 0, 0.5, and 2. Figure 8 shows how $\alpha$ affects the performance achieved by each policy when tested in the five benchmarks, presenting statistics of gains/losses in throughput (see Figure 8) and spectral efficiency (see Figure 8) relative to the OLLA baseline, as well as the average BLER (see Figure 8) and average MCS index selected (see Figure 8), with 95% confidence intervals. Moreover, Figures 9 and 9 further examine how $\alpha$ influences the learned policy by showing the cumulative distribution function (CDF) of the selected MCS index (i.e., the action distribution) and the corresponding BLER, respectively, for benchmark B2.
[Cell throughput statistics vs. $\alpha$.]Refer to caption [Spectral efficiency statistics vs. $\alpha$.]Refer to caption
[Mean BLER vs. $\alpha$.]Refer to caption [Mean MCS index selected vs. $\alpha$.]Refer to caption
Figure 8: Effect of robustness weight $\alpha$ in the reward function on key performance metrics across five benchmark scenarios (B1–B5). Each column of panels represents to a different performance measure evaluated at three robustness weight levels: no penalty ($\alpha=0$), moderate penalty ($\alpha=0.5$), strong penalty ($\alpha=2$). (a) Distribution of percentage gains in downlink cell throughput relative to a baseline policy, summarized as boxplots from multiple Monte Carlo trials (boxes: interquartile range; whiskers: $2\times\text{IQR}$; circles: outliers; diamonds: means). (b) Distribution of percentage gains in spectral efficiency, using the same boxplot conventions as panel (a). (c) Mean BLER with 95% confidence intervals, demonstrating systematic reductions in error rates with increasing retransmission penalization across scenarios. (d) Mean MCS index selected, with 95% confidence intervals, indicating that higher values of $\alpha$ lead to more conservative (lower-rate) transmission modes.
[Action distributions for different $\alpha$.] Refer to caption [BLER distributions for different $\alpha$.]Refer to caption
Figure 9: Effect of robustness weight $\alpha$ in the reward function on the learned RL policy, illustrated through action distributions (MCS index) and BLER for three polices evaluated in benchmark scenario B2 from Figure 8. Increasing $\alpha$ from 0 to 2 encourages more conservative MCS selection, resulting in smaller data payloads better matched to link capacity, thereby reducing BLER distribution.
Setting $\boldsymbol{\alpha=0}$ deactivates the penalty term for failed transmissions (i.e., NACK receptions). This results in a policy that encourages the selection of more aggressive (i.e., higher) MCS indices, as shown in Figure 9, thereby increasing the raw symbol rate. Consequently, the policy learned with $\alpha=0$ tends to maximize spectral efficiency, as illustrated in Figure 8 across all benchmarks, but at the expense of lower throughput, as depicted in Figure 8. Specifically, the policy exploits retransmissions to deliver more information bits than the channel can support in a single transmission, resulting in higher BLER (see Figure 9), which ultimately degrades throughput.
Setting $\boldsymbol{\alpha}$ too high (e.g., $\boldsymbol{\alpha=2}$) has the opposite effect, as it heavily penalizes failed transmissions. The resulting policy promotes very conservative MCS selection, as shown in Figure 9 (in green). Selecting overly low MCS indices leads to transmissions with lower spectral efficiency that carry fewer information bits—but with high reliability. As a result, the policy improves transmission reliability, as evident from the lower BLER distribution in Figure 9, at the cost of reduced throughput and spectral efficiency across all benchmarks, as shown in Figures 8 and 8, respectively.
More precisely, we can identify three clear trends:
  • BLER consistently decreases as $\alpha$ increases, dropping from double-digit percentages at $\alpha=0$ to low single-digit percentages at $\alpha=2$.
  • Average MCS index selection declines with increasing $\alpha$, resulting in lower modulation orders and code rates, and reduced data transmission.
  • Spectral efficiency decreases monotonically with increasing $\alpha$, particularly in high-capacity scenarios such as B4, where it drops from approximately $+40\%$ at $\alpha=0$ to about $+15\%$ at $\alpha=2$.
Thus, increasing $\alpha$ effectively trades off spectral efficiency and data rate for improved reliability.
Setting $\boldsymbol{\alpha}$ to optimize throughput requires a moderate penalty that balances spectral efficiency and reliability. In our evaluations, this balance occurs near $\alpha=0.5$. At this value, the policy reduces retransmissions without overly suppressing MCS aggressiveness, thereby improving throughput. This behavior is observed in most multi-cell scenarios from Table 2—e.g., B2(a/b), B3, B4, and B5—as well as in the interference-free scenario B1, where performance losses at $\alpha=0$ become modest gains at $\alpha=0.5$. Hence, the throughput curve exhibits a concave profile, peaking near $\alpha\approx 0.5$.
Consequently, system designers should select $\alpha$ sufficiently high to mitigate excessive retransmissions—typically around 0.5 in these scenarios—without unduly compromising capacity through overly conservative behavior.
Figure 10: Overview of the proposed architecture. The UE-centric network view (bottom left) is used to construct a graph where nodes represent the serving and neighboring cells. Static and semi-static features (e.g., configuration parameters) are processed by a GNN (e.g., GCN/GAT), implemented as a feedforward architecture with $L$ layers (e.g., GCN or GAT). Dynamic features (e.g., UE measurements) are provided directly to an MLP. Both the GNN and MLP branches are jointly trained via backpropagation, enabling the model to capture both topological cell relations and instantaneous context.
Figure 11: Comparison between MLP-only and GNN-based models (GCN + MLP and GAT + MLP, as per Figure 10) for two versions of benchmark B2(b). Within each plot, the left-most group presents results for 1-site 1-cell version, while the right-most group refers to a denser 3-site 9-cell deployment. Images from left to right show average downlink cell throughput and spectral efficiency gain with respect to OLLA, as well as BLER, respectively.
8.5 Impact of Graph-Based Approach
We next analyze how using graph representations to encode static and semi-static network information—capturing both site/cell attributes and inter-site/cell relationships within a RAN topology—can improve the robustness and transferability of RL models across diverse deployment scenarios. Our findings, presented in Figure 11, indicate that attention-based graphs provide benefits even in relatively small network deployments (e.g., 9 cells), although the gains diminish in very small deployments (e.g., 3 cells), where there is insufficient structural diversity to exploit.
The results in Figure 11 are obtained by encoding the static and semi-static network information outlined in Section 7.3.1, such as cell type, number of antennas, downlink transmit power, and average cell utilization, using a GNN. The overall model architecture, illustrated in Figure 10, consists of a cascade of a GNN followed by an MLP for Q-value estimation. The MLP receives as input both the encoding produced by the GNN and the raw features representing the fast dynamics of link adaptation.
As introduced in Section 4, the input to the GNN consists of attribute vectors $\mathbf{h}_{i}$ for nodes $i$ in a graph representing the local network deployment surrounding a UE, where each node corresponds to either the serving cell or an interfering cell. The full set of node features is denoted by the matrix $\mathbf{H}^{(0)}$, where each row corresponds to the attribute vector of a node, and $\mathbf{H}^{(l)}$ represents the encoding produced at layer $l$ of the GNN.
We consider two GNN-based architectures: GCN + MLP and GAT + MLP. In both cases, the GNN and subsequent MLP are trained end-to-end using backpropagation, minimizing Q-value prediction error following the procedure described in Section 8.1. While we tested all five benchmarks from Table 2, Figure 11 presents results only for benchmark B2(b), as similar conclusions were observed across benchmarks. For this analysis, we also include an extended version of B2 with three sites and nine cells.
We compare these graph-based models with the MLP-only architecture introduced in Section 8.3, which consists of a 6-layer MLP with 256 neurons per layer and no GNN-based encoding. All models are trained with a robustness weight of $\alpha=0.5$ in Equation 2, and we use GNNs with a single layer due to the fully connected graph topology and to reduce computational complexity.
Figure 11 presents results for all approaches in terms of: (a) throughput gain relative to the OLLA baseline, (b) spectral efficiency gain relative to the OLLA baseline, and (c) average experienced BLER. We observe that, while GNN-based and MLP-only models exhibit similar performance in small deployments (e.g., B2(b)), the GAT model achieves approximately 30% higher throughput than both the GNN-based and MLP-only models in the extended 9-cell deployment. This suggests that incorporating attention mechanisms, such as those in the GAT architecture, becomes increasingly beneficial in larger network surrounding, where the topology and interrelations between radio sites/cells become more complex and diverse.
Although our analysis is limited to 9-cell scenarios to reduce simulation time and relies on a constrained set of randomization parameters, the results indicate that graph-based encodings—particularly attention-based ones—may be instrumental for real-world deployments, where cells often operate in highly diverse, irregular, and heterogeneous environments. This is the focus of our ongoing research.
9 CONCLUSIONS AND FUTURE EXTENSIONS
This work explored a critical challenge in the application of deep RL in radio access networks: achieving robust generalization across diverse and dynamic wireless environments. While RL holds promise for automating complex RAN control functions, such as those involved in RRM and network orchestration, its adoption in real-world RANs can be hindered when models trained in specific conditions fail to transfer across the heterogeneous deployment conditions, user behavior, channel dynamics, and so forth.
To address this challenge, we introduced a comprehensive framework to improve generalization in RL-based RAN applications. We began by formally defining RL generalization in the context of RANs, identifying key obstacles such as partial observability and the diversity and non-stationarity of radio environments. We then proposed three core enablers for the development of generalizable RL policies for RAN applications: improved state abstraction tailored to RAN-specific characteristics; diversification of training environments to expose models to a wide range of deployment scenarios and radio environment conditions; and a scalable architecture for distributed learning in RAN to support large-scale data collection across many simulated conditions or real network nodes. Our proposed AI architecture for RAN not only supports sim2sim generalization but also aligns with the structural and operational requirements of real-world RANs, making it suitable for deployment in live networks.
We validated the feasibility and advantages of our approach by designing a novel RL algorithm for link adaptation. We demonstrated, using a high-fidelity system-level simulator compliant with 5G NR systems, that RL algorithms trained with our generalization enablers can outperform fine-tuned models as well as state-of-the-art heuristics when evaluated across diverse and unseen environments. These results highlight the potential of deploying zero-shot generalization in realistic network settings, avoiding frequent retraining.
Overall, this work provides a concrete step toward scalable, generalizable AI solutions for wireless networks. By enabling policy robustness and simplifying lifecycle management, it paves the way for the practical integration of RL solutions into next-generation AI-native RANs.